There are 5 successive compressions of a sound wave train in a distance of 7.2 cm. If
the sound wave can travel 6.8 km in 20 seconds, determine the source frequency (in kHz).
1
Expert's answer
2020-06-17T09:30:26-0400

First, find the wavelength:


"\lambda=\frac{L}{N}=\frac{7.2}{5}=1.44\text{ cm}."

Second, calculate wave velocity:


"v=\frac{d}{t}=\frac{6800}{20}=340\text{ m/s}."

The frequency is


"f=\frac{v}{\lambda}=\frac{340}{0.0144}=23611\text{ Hz},"

or 23.6 kHz.